A population consists of the numbers 2, 4, 5, 9, 10. List all possible sample size of 3 from this population without replacement and determine the mean of each sample.

1
Expert's answer
2022-04-25T08:31:31-0400

The number of possible samples which can be selected without replacement is

"\begin{pmatrix}
 N \\
 n
\end{pmatrix}=\cfrac{N! } {n! \cdot(N-n)! }=\cfrac{5! } {3! \cdot2! }=10."




"\textnormal{Sample\qquad\qquad Mean}\\ 
\ \\
2,4,5\qquad \cfrac{2+4+5}{3}=3\cfrac23\\
2,4,9\qquad \cfrac{2+4+9}{3}=5\\
2,4,10\qquad \cfrac{2+4+10}{3}=5\cfrac13\\
2,5,9\qquad \cfrac{2+5+9}{3}=5\cfrac13\\
2,5,10\qquad \cfrac{2+5+10}{3}=5\cfrac23\\
2,9,10\qquad \cfrac{2+9+10}{3}=7\\
4,5,9\qquad \cfrac{4+5+9}{3}=6\\
4,5,10\qquad \cfrac{4+5+10}{3}=6\cfrac13\\
4,9,10\qquad \cfrac{4+9+10}{3}=7\cfrac23\\
5,9,10\qquad \cfrac{5+9+10}{3}=8"